Other Work

• User Interface and Expert System

– Need to work closely with NiMo operators and engineers on user requirements and rules for expert system

– Nilan and Hariri have experience in previous work for NiMo

• Unit Commitment

– Build on previous work by Dave Koester on shared memory parallel algorithms

– Port to distributed memory, improve efficiencies

– Improve optimization techniques

• Survey of HPC Applications

Possible applications include:

– Customer databases

– Simulation of nuclear power stations

– Natural gas distribution

– Others to be investigated General surveys of this kind have been done by other groups, but have not concentrated on crucial software issues. We will also look at applications important to NiMo.